# Learning Goal-Conditioned Policies from Sub-Optimal Offline Data via Metric Learning

### v2

@@ Abstract
We address the problem of learning optimal behavior from sub-optimal datasets for goal-conditioned offline reinforcement learning. To do so, we propose the use of metric learning to approximate the optimal value function for goal-conditioned offline RL problems under sparse rewards, invertible actions and deterministic transitions. We introduce distance monotonicity, a property for representations to recover optimality and propose an optimization objective that leads to such property. We use the proposed value function to guide the learning of a policy in an actor-critic fashion, a method we name MetricRL. Experimentally, we show that our method estimates optimal behaviors from severely sub-optimal offline datasets without suffering from out-of-distribution estimation errors. We demonstrate that MetricRL consistently outperforms prior state-of-the-art goal-conditioned RL methods in learning optimal policies from sub-optimal offline datasets.
@@ 1 Introduction
Figure 1: Average reward on Minigrid DoorKey [1] as a function of the expected reward present in the offline dataset. We contribute MetricRL (red line), a novel goal-conditioned offline RL agent able to learn near-optimal behavior from severely sub-optimal datasets.
@@
Effective decision-making is an integral part of intelligent behavior. To achieve this, learning-based control methods have proven to be a viable option in complex scenarios [2, 3, 4, 5]. In particular, reinforcement learning (RL) allows learning near-optimal behavior through trial-and-error [6]. However, the online reinforcement learning framework generally requires slow, expensive (and potentially dangerous) online interactions with the environment.
@@
Offline RL, on the other hand, formalizes the learning of optimal behaviors from a static dataset [7]. This approach offers many advantages over its online counterpart, such as the ability to leverage large-scale datasets to learn complex behavior [8, 9] without the need of re-collecting data [10, 11]. Because of the inability to access the environment, in offline RL it is assumed that the dataset already includes suitable information to perform the given task. This data, however, may often be collected by sub-optimal agents. In this work, we address the question of how to learn different (and better) behaviors from the one observed in the dataset, regardless of its optimality. We focus on the challenge of learning near-optimal behavior from severely sub-optimal datasets, such as the ones collected by a random policy. As we empirically demonstrate in Section 4, and highlight in Figure 1, current offline RL methods [12, 13, 14, 15] struggle significantly under these conditions.
@@
We focus on sparse-reward goal-conditioned offline reinforcement learning which aims at learning optimal behavior to reach multiple goals within the same environment. Motivated by recent work in quasimetric learning [13], we explore the use of metric learning to estimate policies from sub-optimal offline data [16, 17]. Differently from previous work, by assuming invertible actions, we can naturally exploit symmetry as a bias in the learning process of our agent. Endowing the agent with this bias allows for learning optimal behaviors in severely sub-optimal data conditions: consider the undirected graph implicitly defined by the offline dataset where states (the nodes) are connected by actions (the edges). The core idea of the proposed method is to exploit symmetries of metric spaces to learn an embedding of this graph such that Euclidean distances are related to the value function. In particular, we introduce the notion of distance monotonicity as a relaxation of isometries of this embedding and show that preserving this property provably allows for the recovery of an optimal policy, regardless of the quality of the dataset used for training. We call our method MetricRL.
@@
We evaluate MetricRL across a wide range of literature-standard goal-conditioned reinforcement learning tasks. We show how MetricRL outperforms prior goal-conditioned offline reinforcement learning methods in learning near-optimal behavior from severely sub-optimal datasets. Additionally, we show how MetricRL easily scales to high-dimensional observations.
@@
  • MetricRL: We propose a novel representation to exploit symmetries in goal-conditioned offline reinforcement learning.
  • Distance Monotonicity: We define a new property of these representations. We show that preserving such property provably leads to policy optimality.
@@ 2 Preliminaries and Assumptions
Goal-Reaching Reinforcement Learning: We consider standard Markov Decision Processes (MDPs) for goal-reaching tasks, ${\mathcal{M}=(S,A,T,r,\gamma)}$, where $S$ is the state space, $A$ is the action space, $T:S\times A\to S$ is a deterministic transition function (a common assumption in recent offline RL methods [14, 18, 13]), $r:S\times A\to\mathbb{R}$ is a goal-conditioned, sparse reward, i.e., $r(s,a)\neq 0$ iff $T(s,a)=s_{g}$ (where $s_{g}$ is the goal-state), and $\gamma\in[0,1)$ is a discount factor on the future rewards of the agent. We additionally define the goal states to be absorbing and consider the process terminated once these are reached.
@@
The goal of the process is to find an optimal policy ${\pi^{*}(a\mid s,s_{g})}$ that, given goal state $s_{g}$, maximizes the cumulative discounted reward of the agent for any possible starting state. To evaluate the optimality of the policy we resort to the goal-conditioned value function $V^{\pi}(s,s_{g})$, defined as the expected discounted cumulative reward starting in a particular state and acting according to a policy ${\pi:{V^{\pi}(s,s_{g})=\operatorname{\mathbb{E}}_{\pi}\left[\sum_{t}\gamma^{t}r_{t}|s_{0}=s,a=\pi(s,s_{g})\right]}\quad\forall s\in S}$. The value function associated with the optimal policy is referred to as the optimal value function $V^{*}=V^{\pi^{*}}$ and, as such, is always greater or equal to any other value functions, i.e. $V^{*}(s,s_{g})\geq V^{\pi}(s,s_{g})\quad\forall s,\pi$.
@@
Offline Reinforcement Learning: In the offline RL setting, we assume we have access to a dataset $D$ of interactions with the environment described by the MDP and collected by some unknown policy $\pi_{\beta}$, i.e. ${D_{\pi_{\beta}}=\{(s,a\sim\pi_{\beta}(a\mid s),r,s^{\prime}=T(s,a))\}}$. As discussed in prior work [13, 19], the dataset $D$ implicitly defines a graph ${G=(S,E)}$ where the nodes correspond to the states and the edges to the actions of the agent. Here we assume that this graph only has one connected component, i.e., any two states in the dataset are connected through a path on the graph. Furthermore, we assume that there always exists an inverse action, i.e., ${\exists a^{\prime}\in A:T(s^{\prime},a^{\prime})=s\quad\forall(s,a,s^{\prime}=T(s,a))}$. Note that $a^{\prime}$ doesn’t necessarily need to correspond to the actual opposite action and it can be any viable action. By endowing the graph $G$ with a metric $d_{G}$, we can define a corresponding finite metric space $(G,d_{G})$. For the tuple to be a valid metric space we need to define the metric $d_{G}:S\times S\to\mathbb{R}$ to respect the axioms of a metric space [20]. In particular, we can define the distance between any two states to be the number of edges on the shortest path connecting them (geodesic distance).
@@ 3 Method
Figure 2: We explore a form of symmetry in representation learning for goal-condition offline reinforcement learning: we learn a metric space in which Euclidean distances between the representation of states ($z,z^{\prime},z^{\prime\prime}$) are related to the value function of the agent. We call our approach MetricRL. In the Minigrid Doorkey environment, moving greedily to adjacent states translates to the optimal policy (red line) to reach the goal (in green).
In this work we focus on learning near-optimal goal-conditioned behavior from sub-optimal offline data. In these conditions, two main problems arise. The first is to learn an optimal behavior without depending on the quality of the distribution of the data used. The second is to avoid the out-of-distribution shift commonly faced in offline reinforcement learning [7]. We propose to address this using a metric learning approach to estimate the optimal value function. To do so, we start by defining distance monotonicity (Section 3.1), a novel property on representations needed to recast the problem of optimal value function estimation into metrics. We propose a loss function to learn maps that respect such property and show how to build an approximation of the value function using distances in this learned representation (Section 3.2). Finally, we define an actor-critic method to learn policies over this approximation and formally prove their optimality (Section 3.3). We call our approach MetricRL.
@@ 3.1 Distance Monotonicity
Consider a continuous map between the state space of an MDP for goal-reaching tasks and a latent representation space: $\phi:S\to Z\subseteq\mathbb{R}^{n}$. We can equip this latent vector space with a Euclidean norm to obtain a Euclidean metric space $(Z,\left\lVert.\right\rVert_{2})$. We say $\phi$ is isometric if relative distances in the original state space $d_{S}$ and the latent metric space $d_{Z}$ are preserved, i.e., ${d_{Z}\left(\phi(s),\phi(s^{\prime})\right)=d_{S}(s,s^{\prime}),\,\forall s,s^{\prime}\in S}$. In fact, if $\phi$ is isometric then the value function can be defined as simply the norm of the distance between the current state and the goal state A.2, i.e., ${V^{*}(s,s_{g})=\gamma^{d_{Z}(\phi(s),\phi(s^{\prime}))}r_{g}}$ where $r_{g}$ represents the reward at the goal and the expectation has been dropped as we assume deterministic transitions.
However, estimating an isometry between these two metric spaces is known to be not always possible [21, 22]. To overcome such issue, we consider a relaxation of isometries between metric spaces. Given two metric spaces $(S,d_{S})$ and $(Z,d_{Z})$ and the corresponding map $\phi$ between them, we can define the following property of $\phi$:
@@
We say $\phi$ is distance monotonic (DM) if for all $s_{1},s_{2},s_{3}\in S$, ${d_{S}(s_{1},s_{3})\leq d_{S}(s_{2},s_{3})}$ implies $d_{Z}(z_{1},z_{3})\leq d_{Z}(z_{2},z_{3})$.
A similar property has been described in [23]. There, however, it is imposed on a learned representation by means of explicit constrain in the optimization procedure. Instead, we propose to parameterize the map $\phi_{\theta}$ and learn it by minimizing the following objective on the dataset $D$:
$\mathcal{L}_{\theta}(D)=\operatorname{\mathbb{E}}_{D}\left[(\left\lVert\phi_{\theta}(s^{\prime})-\phi_{\theta}(s)\right\rVert_{2}-1)^{2}-\lambda\left\lVert\phi_{\theta}(s^{\prime\prime})-\phi_{\theta}(s)\right\rVert_{2}\right],$ (1)
Figure 3: Optimizing Equation  1 increases the ratio of distance monotonic triplets (blue curve) on Maze2D (Large). Distance monotonicity is also correlated with an increase in the average return of the agent (orange curve).
@@
The minimization of Equation 1 enforces the learning of a distance monotonic map: in Figure 3 we highlight that the ratio of distance monotonic triplets significantly increases as a function of the training of the map. In Appendix A.3 we provide a more formal intuition of this relationship as well as additional details on the evaluation of the distance monotonicity ratio.
@@ 3.2 Value Function Approximation
However, distance monotonicity is enough to retain relative distances between states. Figure 4 shows the estimated value function of different algorithms for a maze-like problem, using offline datasets collected with policies of different values (low, medium and high). In such problems the value function should retain the topology of the maze and estimate the value of each position in the maze based on the distance to the goal within the maze.
Figure 4 highlights that a distance monotonic representation (MetricRL) is the only value function able to correctly estimate the low value of the bottom left corner of the maze when the goal is on the other branch of the maze, regardless of the quality of the policy used to collect the offline dataset. Equations 1 and 2 allow us to abstract the estimate of the value function from the distribution of the policy that collected the dataset. Note that, for more complex topologies (e.g., loop on the top-right corner), distance monotonicity is not equivalent to isometries. However, as we show in the next subsection, our approximation still allows us to estimate a provably optimal policy.
Figure 4: Estimated value function for different methods in Maze2D large using datasets collected using policies of different values (rows). Values are normalized.
@@ 3.3 MetricRL
Enforcing distance monotonicity in the latent representation allows us to learn an approximation of the true value function, not fully recover it. However, when distance monotonicity is preserved a greedy policy built on this value function is equivalent to the optimal policy. Define the greedy policy according to the value function $V$ as ${\pi_{g}^{V}(s)=\operatorname*{arg\,max}_{a}V(T(s,a))}$. We have the following:
@@
$\pi_{g}^{\tilde{V}}(s)=\pi_{g}^{V^{*}}(s)\quad\forall s\in S$
@@
A proof can be found in Appendix A.4.
To learn such policy we take a three-stage approach, which we call MetricRL: (i) we learn a distance monotonic map, using the loss function of Equation 1; (ii) we define the value function using Equation 2; (iii) we employ an actor-critic approach [24] to learn a policy using policy gradient loss function:
$\mathcal{L}_{\pi}=\operatorname{\mathbb{E}}_{s,a\sim D}\left[(\tilde{V}(s^{\prime})-\tilde{V}(s))\log(\pi(a\mid s))\right].$ (3)
This actor-critic approach (similar to [25]) is particularly suitable for offline RL: actions are sampled from the dataset which guarantees us to consider only in-distribution actions, solving the classic offline RL issue of out-of-distribution transitions [7]. Moreover, using a value function from a distance monotonic representation ensures the term ${\tilde{V}(s^{\prime})-\tilde{V}(s)}$ to be positive only for actions that bring the agent closer to the goal.

### v3

# Goal-Conditioned Reinforcement Learning from Sub-Optimal Data on Metric Spaces

@@ Abstract
We study the problem of learning optimal behavior from sub-optimal datasets for goal-conditioned offline reinforcement learning under sparse rewards, invertible actions and deterministic transitions. To mitigate the effects of distribution shift, we propose MetricRL, a method that combines metric learning for value function approximation with weighted imitation learning for policy estimation. MetricRL avoids conservative or behavior-cloning constraints, enabling effective learning even in severely sub-optimal regimes. We introduce distance monotonicity as a key property linking metric representations to optimality and design an objective that explicitly promotes it. Empirically, MetricRL consistently outperforms prior state-of-the-art goal-conditioned RL methods in recovering near-optimal behavior from sub-optimal offline data.
@@ 1 Introduction
Figure 1: Average reward on Minigrid DoorKey (Chevalier-Boisvert et al., 2023) as a function of the expected reward present in the offline dataset. We contribute MetricRL (red line), a novel goal-conditioned offline RL agent able to learn near-optimal behavior from severely sub-optimal datasets.
@@
Effective decision-making is an integral part of intelligent behavior. To achieve this, learning-based control methods have proven to be a viable option in complex scenarios (Andrychowicz et al., 2020; Silver et al., 2017; Mnih et al., 2013; Peters and Schaal, 2008). In particular, reinforcement learning (RL) allows learning near-optimal behavior through trial-and-error (Sutton and Barto, 2018). However, the online reinforcement learning framework generally requires slow, expensive (and potentially dangerous) online interactions with the environment.
@@
Offline RL, on the other hand, formalizes the learning of optimal behaviors from a static dataset (Levine et al., 2020). This approach offers many advantages over its online counterpart, such as the ability to leverage large-scale datasets to learn complex behavior (Walke et al., 2023; Dasari et al., 2020) without the need of re-collecting data (Shi et al., 2021; Gürtler et al., 2023). Because of the inability to access the environment, in offline RL it is assumed that the dataset already includes suitable information to perform the given task. This data, however, may often be collected by sub-optimal agents. In this work, we address the question of how to learn different (and better) behaviors from the one observed in the dataset, regardless of its optimality. We focus on the challenge of learning near-optimal behavior from severely sub-optimal datasets, such as the ones collected by a random policy. We empirically demonstrate in Section 4, and highlight in Figure 1, how under these conditions current offline RL methods (Eysenbach et al., 2022; Wang et al., 2023; Ma et al., 2022b; Kostrikov et al., 2021) struggle significantly.
@@
We focus on sparse-reward goal-conditioned offline reinforcement learning, which aims at learning optimal behavior to reach multiple goals within the same environment. Motivated by recent work in metric (Park et al., 2024; 2023b) and quasimetric (Eysenbach et al., 2022; Wang et al., 2023) learning for RL, we study an approximation of the optimal value function in severely sub-optimal data conditions without explicitly relying on the Bellman operator, (Bellman, 1954). The core idea is to learn an embedding of the state space such that distances correspond to the minimum number of actions needed to reach one state from the other. We introduce the notion of distance monotonicity (DM) as a relaxation of isometric embeddings and show that this property is sufficient to provably recover an optimal actor-critic policy. In particular, we propose to pair this approximation of the value function with a weighted imitation learning actor, a method we refer to as MetricRL, to recover an optimal policy regardless of the dataset quality. This is made possible by avoiding the out-of-distribution issue caused by the $\max$ operator used in dynamic programming solutions, without requiring any additional conservative (e.g., as in CQL (Kumar et al., 2020)) or behavioral (e.g., as in BCQ (Fujimoto et al., 2019), BEAR (Kumar et al., 2019)) regularization.
@@
In this work, we define the learned embedding as a Euclidean metric space. This effectively limits the applicability of the model to reversible processes (Steccanella and Jonsson, 2022). On the other hand, we empirically show that this induced bias can ease the learning process in offline RL for many established environments. We evaluate MetricRL across a wide range of literature-standard goal-conditioned reinforcement learning tasks. We show how MetricRL outperforms prior goal-conditioned offline reinforcement learning methods in learning near-optimal behavior from severely sub-optimal datasets. Additionally, we show how MetricRL easily scales to high-dimensional observations.
@@
  • MetricRL: We propose a novel method that exploits symmetries in latent representation spaces for reversible goal-conditioned offline reinforcement learning.
  • Distance Monotonicity: We define a new property of these latent representation spaces. We show that, under invertible actions, preserving such a property provably leads to policy optimality.
@@ 2 Preliminaries and Assumptions
Goal-Reaching Reinforcement Learning: We consider standard Markov Decision Processes (MDPs) for goal-reaching tasks, $\mathcal{M}=(S,A,T,r,\gamma)$}, where $S$ is the state space, $A$ is the action space, $T:S\times A\to S$ is a deterministic transition function (a common assumption in recent offline RL methods (Ma et al., 2022b; Park et al., 2023a; Wang et al., 2023)), $r:S\times A\to\mathbb{R}$ is a goal-conditioned, sparse reward, i.e., $r(s,a)\neq 0$ iff $T(s,a)=s_{g}$ (where $s_{g}$ is the goal-state), and $\gamma\in[0,1)$ is a discount factor on the future rewards of the agent. We additionally define the goal states to be absorbing and consider the process terminated once these are reached.
@@
The goal of the process is to find an optimal policy $\pi^{*}(a\mid s,s_{g})$} that, given goal state $s_{g}$, maximizes the cumulative discounted reward of the agent for any possible starting state. To evaluate the optimality of the policy we resort to the goal-conditioned value function $V^{\pi}(s,s_{g})$, defined as the expected discounted cumulative reward starting in a particular state and acting according to a policy $\pi:{V^{\pi}(s,s_{g})=\mathbb{E}_{\pi}\left[\sum_{t}\gamma^{t}r_{t}|s_{0}=s,a=\pi(s,s_{g})\right]}\quad\forall s\in S$}. The value function associated with the optimal policy is referred to as the optimal value function $V^{*}=V^{\pi^{*}}$ and, as such, is always greater or equal to any other value functions, i.e., $V^{*}(s,s_{g})\geq V^{\pi}(s,s_{g})\quad\forall s,\pi$.
@@
Offline Reinforcement Learning: In the offline RL setting, we assume we have access to a dataset $D$ of interactions with the environment described by the MDP and collected by some unknown policy $\pi_{\beta}$, i.e., $D_{\pi_{\beta}}=\{(s,a\sim\pi_{\beta}(a\mid s),r,s^{\prime}=T(s,a))\}$}. A major issue in Offline RL stems from the way dynamic programming methods estimate the optimal value function of the MDP. Most of the current algorithms rely on temporal difference techniques to learn the critic function, e.g., DQN (Mnih et al., 2013), CQL (Kumar et al., 2020), BEAR (Kumar et al., 2019), BCQ (Fujimoto et al., 2019). However, the $\max$ operator used to estimate the target value has been shown to result in an overestimation of the expected return, Levine et al. (2020). While this is not an issue in Online RL, as the agent has the possibility of exploring overestimated states, it results in catastrophic effects in Offline RL. In this paper, we rely on an alternative technique for learning the critic. As discussed in prior work (Wang et al., 2023; Yang et al., 2020), the dataset $D$ implicitly defines a graph G=(S,A){G=(S,A)} where the nodes correspond to the states and the edges to the actions of the agent. Here, we assume that this graph only has one connected component, i.e., any two states in the dataset are connected through a path on the graph. Furthermore, we assume that there always exists an inverse action, i.e., $\exists a^{\prime}\in A:T(s^{\prime},a^{\prime})=s\quad\forall(s,a,s^{\prime}=T(s,a))$}. Note that $a^{\prime}$ doesn’t necessarily need to correspond to the actual opposite action and it can be any viable action. By endowing the graph $G$ with a metric $d_{S}$, we can define a corresponding finite metric space $(G,d_{S})$. For the tuple to be a valid metric space we need to define the metric $d_{S}:S\times S\to\mathbb{R}$ to respect the axioms of a metric space (Burago et al., 2001). In particular, we can define the distance between any two states to be the number of edges on the shortest path connecting them (geodesic distance).
@@ 3 Method
Figure 2: We explore a form of symmetry in representation learning for goal-conditioned offline reinforcement learning: we learn a metric space in which Euclidean distances between the representation of states ($z,z^{\prime},z^{\prime\prime}$) are related to the value function of the agent. We call our approach MetricRL. In the Minigrid Doorkey environment, moving greedily to adjacent states translates to the optimal policy (red line) to reach the goal (in green). Our objective is to preserve the local structure of adjacent representations (orange arrows, left) while maximizing the separation between non-adjacent ones (blue arrows, left).
In this work we focus on learning near-optimal goal-conditioned behavior from sub-optimal offline data. In these conditions, two main problems arise. The first is to learn an optimal behavior without depending on the quality of the distribution of the data used. The second is to avoid the out-of-distribution shift commonly faced in offline reinforcement learning (Levine et al., 2020). We propose to address both using a metric learning approach to estimate the optimal value function. To do so, we start by defining distance monotonicity (Section 3.1), a novel property on representations needed to recast the problem of optimal value function estimation into metrics. We propose a loss function to learn maps that respect such a property and show how to build an approximation of the value function using distances in this learned representation (Section 3.2). Finally, we define an actor-critic method to learn policies over this approximation and formally prove their optimality (Section 3.3). We call our approach MetricRL.
@@ 3.1 Distance Monotonicity
Consider a continuous map between the state space of an MDP for goal-reaching tasks and a latent representation space: $\phi:S\to Z\subseteq\mathbb{R}^{n}$. We can equip this latent vector space with a Euclidean norm to obtain a Euclidean metric space $(Z,\left\lVert.\right\rVert_{2})$. Here we consider distances in the original space, $d_{S}$ as the minimum number of actions an optimal policy needs to reach one state from the other. We say $\phi$ is isometric if relative distances in the original state space $d_{S}$ and the latent metric space $d_{Z}$ are preserved, i.e., $d_{Z}\left(\phi(s),\phi(s^{\prime})\right)=d_{S}(s,s^{\prime}),\,\forall s,s^{\prime}\in S$}. In fact, if $\phi$ is isometric then the value function can be defined as simply the norm of the distance between the current state and the goal state, i.e., $V^{*}(s,s_{g})=\gamma^{d_{Z}(\phi(s),\phi(s^{\prime}))}r_{g}$} where $r_{g}$ represents the reward at the goal and the expectation has been dropped as we assume deterministic transitions. We present an extended discussion of this observation in Appendix A.1.
However, estimating an isometry between these two metric spaces is known to be not always possible (Bourgain, 1985; Matoušek, 2002). To overcome such an issue, we consider a relaxation of isometries between metric spaces. Given two metric spaces $(S,d_{S})$ and $(Z,d_{Z})$ (e.g., $d_{S}$ the geodesics in the state space and $d_{Z}$ a simple $\ell_{2}$-norm) and the corresponding map $\phi$ between them, we can define the following property of $\phi$:
@@
We say $\phi$ is distance monotonic (DM) if for all $s_{1},s_{2},s_{3}\in S$, the following holds:
$d_{S}(s_{1},s_{3})<d_{S}(s_{2},s_{3})\implies d_{Z}(\phi(s_{1}),\phi(s_{3}))<d_{Z}(\phi(s_{2}),\phi(s_{3})).$
We propose to parameterize the map $\phi_{\theta}$ and learn it by minimizing the following objective on the dataset $D$:
$\mathcal{L}_{\theta}(D)=\mathbb{E}_{D}\left[(\left\lVert\phi_{\theta}(s^{\prime})-\phi_{\theta}(s)\right\rVert_{2}-1)^{2}-\lambda\left\lVert\phi_{\theta}(s^{\prime\prime})-\phi_{\theta}(s)\right\rVert_{2}\right],$ (1)
Figure 3: Optimizing Equation 1 increases the ratio of distance monotonic triplets (blue curve) on Maze2D (Large). Distance monotonicity is also correlated with an increase in the average return of the agent (orange curve).
@@
The minimization of Equation 1 enforces the learning of a distance monotonic map: in Figure 3 we highlight that the ratio of distance monotonic triplets significantly increases as a function of the training of the map. This can be approximated by discretizing the state space, building an $\epsilon$-graph and the resulting distances in the state space as geodesics on the graph, comparing these distances with distances in the learned representation $Z$. In Appendix A.2 we provide a more formal intuition of this relationship as well as additional details on the evaluation of the distance monotonicity ratio.
@@ 3.2 Value Function Approximation
However, distance monotonicity is enough to retain relative distances between states. Figure 4 shows the estimated value function of different algorithms for a maze-like problem, using offline datasets collected with a random policy. In such problems the value function should retain the topology of the maze and estimate the value of each position in the maze based on the distance to the goal within the maze.
Figure 4 (red) highlights that a distance monotonic representation (MetricRL) is the only value function able to correctly estimate the low value of the bottom left corner of the maze when the goal is on the other branch of the maze. Equations 1 and 2 allow us to abstract the estimate of the value function from the distribution of the policy that collected the dataset, as we highlight in Figure 11 for policies of different values. Note that, for more complex topologies (Figure 4, orange), distance monotonicity is not equivalent to isometries. However, as we show in the next subsection, our approximation still allows us to estimate a provably optimal policy.
Figure 4: Estimated value function for different methods using a dataset collected from a random policy in Maze2D Large. We highlight (in red) how MetricRL is the only method able to correctly assign low values to states that are close in the Euclidean space but not in terms of distance to the goal. Additionally, we highlight (in orange) that our proposed distance monotonicity in complex topologies is not equivalent to isometries, yet we are still able to recover provably optimal policies (as we discuss in Section 3.3). All values are normalized.
@@ 3.3 MetricRL
Enforcing distance monotonicity in the latent representation allows us to learn an approximation of the true value function, not fully recover it. However, when distance monotonicity is preserved, a greedy policy built on this value function is optimal given some assumptions. Define the greedy policy according to the value function $V$ as $\pi_{g}^{V}(s)=\operatorname*{arg\,max}_{a}V(T(s,a))$}, note that in general the argmax results in a set of possible actions. We have the following:
@@
$\pi_{g}^{\tilde{V}}(s)\subseteq\pi_{g}^{V^{*}}(s),\quad\forall s\in S$
@@
Intuitively, Theorem 3.2 states that every policy greedy on the value function defined in Equation 2 is optimal. A proof can be found in Appendix A.3. When the actions space is too large or continuous, we pair the learned representation with a parametric policy.
To learn such policy we take a three-stage approach, which we call MetricRL: (i) we learn a distance monotonic map, using the loss function of Equation 1; (ii) we define the value function using Equation 2; (iii) we estimate a policy using the following loss function:
$\mathcal{L}_{\pi}=\mathbb{E}_{s,a\sim D}\left[-(\tilde{V}(s^{\prime})-\tilde{V}(s))\log(\pi(a\mid s))\right].$ (3)
This optimization procedure for the policy is akin to the family of weighted maximum likelihood Nair et al. (2020) or weighted imitation learning Wang et al. (2018). A key difference is that the advantage is not exponentiated. Empirically, this didn’t have particular effects on the learning of the policy and allowed the removal of the additional hyper-parameter controlling the temperature of the exponential, thus simplifying the training procedure. Using a value function from a distance monotonic representation ensures the term $\tilde{V}(s^{\prime})-\tilde{V}(s)$} is positive only for actions that bring the agent closer to the goal. This approach for the policy update is particularly suitable for offline RL: actions are sampled from the dataset which guarantees us to consider only in-distribution actions. Moreover, using the proposed distance monotonic representation instead of Temporal Difference methods for the critic solves the classic offline RL issue of out-of-distribution transitions (Levine et al., 2020). We provide a pseudocode of our approach in Appendix A.5.